Maisha went to Berkeley Bowl and spent $9.20 on 6 mangos and 5 oranges. Talida stopped at Berkeley Bowl on the same day and paid $7.60 for 3 mangos 10 oranges. Show how to use equations to figure out the prices for mangos and oranges that day at Berkeley Bowl.
Answer by addingup(3677)

6m+5o= 9.20
and
3m+10o= 7.60 Multiply all sides times 2:
6m+20o= 15.20 Now subtract this from the first equation:
------------------
6m+5o= 9.20
- 6m+20o= 15.20
--------------------
-15o= -6 Divide both sides by -15:
o= 0.40 each. And the mangos?
Maisha:
5 oranges at 0.40= 2.00
6 mangos: 9.20-2= 7.20; 7.20/6= 1.20 each If our answer is correct, we should get the same for Talida.
Talida:
10 oranges at 0.40= 4.00
3 mangos: 7.60-4= 3.60/3= 1.20 each
As you see, we have the correct answer. They both paid 0.40 per orange and 1.20 per mango.